The area of the wall you are going to paint is 50 feet square. If the width is half the height what are the dimensions of the wall?
Answer by ewatrrr(24785)


Hi
width is half the height
Let x and x/2 represent height and width respectively
Question states***
x(x/2) = 50 ft^2
solving for x
x^2/2 = 50 ft^2
x^2 = 100 ft^2
x = ± 10 |Tossing out negative solution for length
x = 10ft, the height of the wall. the width is 5ft ((1/2)*10ft)